OpenStudy (anonymous):

2√3 cos (theta/2)=-3

OpenStudy (anonymous):

2√3 cos (theta/2)=-3 cos (theta/2)=-3/(2√3) cos (theta/2)= -√3/2

OpenStudy (anonymous):

cos (theta/2) = -cos60 cos (theta/2) = cos(180-120) cos (theta/2) = cos120 theta /2 = 120 theta =240

OpenStudy (anonymous):

thanks so much!